ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

0. Plenary Lectures

1991 MS Classification: Sarnak, Peter, Princeton University, USA Zeta and L-functions

The theory of zeta and L-functions is at the center of a number of recent developments in number theory. We will review some of these developments ranging from the Riemann Zeta function to standard L-functions for automorphic forms on the general linear group. Applications to classical problems in number theory and related areas will be highlighted.

ICM 1998 Berlin · Plenary and Invited Lectures